OpenStudy (anonymous):

please help, points U, W, X, and Y are collinear. How many line segments are determined?

OpenStudy (anonymous):

collinear = falling on same line therefore you have many options do you have more details l

OpenStudy (anonymous):

it is based on scenarios example u-w-x-y u-x-w-y and so on each conditions gives you a set of line segments

OpenStudy (australopithecus):

\[\frac{1}{4}\frac{x}{1} - 5 = -2\] \[\frac{1}{4}\frac{x}{1} - 5 +5 = -2 + 5\] \[\frac{1}{4}\frac{x}{1} - 0 = 3\] \[\frac{1}{4}\frac{x}{1} = 3\] \[\frac{1(x)}{4(1)} = 3\] \[\frac{(x)}{4} = 3\] \[\frac{(x)}{4}*\frac{4}{1} = 3*4\] \[\frac{(x)(4)}{4(1)} = 12\] \[\frac{(x)(1)}{1} = 12\] \[\frac{(x)}{1} = 12\] x = 12

OpenStudy (australopithecus):

you can express any number as a fraction by using 1 as the denominator

OpenStudy (australopithecus):

and fraction multiplication is easy its just numerator*numerator & denominator*denominator

OpenStudy (anonymous):

ok thanks so there is no way of doing this using purely geometric method?

OpenStudy (australopithecus):

wow I posted this in the wrong question lol damit

OpenStudy (australopithecus):

my track pack has a mind of its own :( and clicks things randomly

OpenStudy (australopithecus):

apple really sucks in the quality department never buying a macbook again

OpenStudy (anonymous):

its ok, could you still help me solve it though please?

OpenStudy (australopithecus):

sorry not familiar with this topic of mathematics

OpenStudy (anonymous):

there is indeed a simple way of doing it lets say you have u -x-y-w then you will have ux, uy, uw, xy, xw, yw as segments think on same lines for all cases for u lets say you have x-y-w-u then you willhave xy, yw, wu, xw, you xu as options thats the reason, i asked for conditions of collinearlity

OpenStudy (anonymous):

ok thanks so you then have 6 line segments?

OpenStudy (anonymous):

many because we can also have y-x-w-u

OpenStudy (anonymous):

24 line segments for y-x-w-u

OpenStudy (anonymous):

similarly for other conditions